OpenStudy (anonymous):

medal!! George tells you that when variables are in the denominator, the equation becomes unsolvable. "There is a value for x that makes the denominator zero, and you can't divide by zero," George explains. Using complete sentences, demonstrate to George how the equation is still solvable.

OpenStudy (anonymous):

@DanJS @jim_thompson5910

OpenStudy (danjs):

GEORGE!!

OpenStudy (anonymous):

huh? lol

OpenStudy (danjs):

The equation is still valid over its domain. For instance, Y= 1/x The domain is All X, such that X does not equal Zero.

OpenStudy (danjs):

Having X in the denominator doesn't make the thing unsolvable, it just imposes restrictions on the domain.

OpenStudy (anonymous):

ooo gotcha.

OpenStudy (danjs):

Graph y = 1/x and look what happens near x=0